• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name L.Pentchev

    Log entry time 23:58:17 on October11,2006

    Entry number 186873

    keyword=Swing shift summary

    Shifters: L.Pentchev, O.Glamazdin

    16:00 starting shift during RUN 1232 4He production, 80uA

    16:09-16:44 run 1233 1.5M

    16:45 found that the right NMR was frozen for less than 5min,
    just before the end of run 1233, rebooted iocha16 and Right dipole NMR, cycled Q2 and Q3, restored angles, lost about 30min

    17:16-17:49 run 1235 (1234 junk) 1.5M
    17:50-18:24 run 1236 1.5M
    18:25-19:02 run 1237 1.5M
    19:02-19:34 run 1238 1.5M
    19:34-20:04 run 1239 1.5M
    20:04-20:40 run 1240 1.5M

    20:40 MCC called, the is beam off for measurements, 20:56 beam is back

    20:57-21:29 run 1241 1.5M
    21:30-22:01 run 1242 1.5M

    22:00 Steffen found that the right dipole momentum starting from run 1240 is frozen, the field monitor for R/L fields was frozen, not showing constant value for the R field as before (at 16:45), that's why we didn't recognize the problem
    After rebooting iocha16 and the Right dipole NMR the problem is solved, it took 15 min.

    22:16-22:54 run 1243 1.5M
    22:54-23:30 run 1244 1.5M
    23:31 run 1245 started